RATE CONSTANTS AND DEUTERIUM KINETIC ISOTOPE EFFECTS FOR METHOXY RADICAL RE-ACTING WITH NO_{2} AND O_{2}

\underline{J. CHAI}, H. HU, T. S. DIBBLE, Department of Chemistry, College of Environmental Science and Forestry, State University of New York, Syracuse, New York 13210; G. S. TYNDALL, J. J. OR-LANDO, Atmospheric Chemistry Division, National Center for Atmospheric Research, Boulder, Colorado 80305.

Alkoxy radicals (RO) are important intermediates in the photooxidation of volatile organic compounds due to their great impacts on ozone formation and gas-particle partitioning of stable oxidation products. Methoxy radical (CH_{3}O) is the prototype for all alkoxy radicals. The absolute rate constants k_{N}_{O}_{2}(T) for reaction of CH_{3}O and CD_{3}Owith NO_{2} have been measured using laser flash photolysis to generate radicals and laser-induced fluorescence for time-resolved detection. The pressure and temperature dependence for k_{N}_{O}_{2} have been observed over the range 30-700 Torr and 250-335 K. This will be the first direct measurement of k_{N}_{O}_{2} for CH_{3}O near ambient pressure and the first ever for CD_{3}O.The relative rate constants k_{N}_{O}_{2}/k_{O}_{2}(T) have been measured in an environmental chamber with FTIR detection. This combination enables the first determination of k_{O}_{2} (T) for CH_{3}O and CD_{3}O for T < 298 K. The results will also help validate theoretical methods for studying alkoxy + O_{2} reactions, which are challenging for quantum chemistry.